OpenStudy (anonymous):

The volume of a cylinder with height h and radius r is given by the formula below. V = Πr 2 h Complete the equation below by writing an equation equivalent to height, h, of a cylinder.

OpenStudy (anonymous):

\(V=\pi r^2 h\), solving for h divide both side by \(\pi r^2 h\), and tell me what you get

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

\(\LARGE \frac{V}{\pi r^2}=\frac{\pi r^2 h}{\pi r^2}\), \(\pi r^2\) , will cancel on the right side so you will have?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what did you get?